Solve the quadratic equation: x^2-2x-4=-x^2+7x+1
Answer by nerdybill(7384)

x^2-2x-4=-x^2+7x+1
2x^2-2x-4=7x+1
2x^2-9x-4=1
2x^2-9x-5= 0
factoring the left side:
(2x+1)(x-5)= 0
x = {-0.5, 5}